PROBLEM I. Suppose that there are 6 price-taker milk farms in a town They produce milk of the same quality, and each has a cost function C() 0.04q2. The market demand function for milk in this town is given by Q(p)- 1600 325p. 1. The market supply function for milk in this town is (a) Q(p)1600 (b) Qs(p) = 75p (c) Q(p) = 125p (d) Q(p 1200 (e) Q(p)=25p 325p. 400p Q2. What is the competitive equilibrium price of milk in this town? (a) $3.00 (b) $3.50 (c) $4.25 (d) $2.75 (e) $4.00
Answer #1

(Q1) (b)

Individual firm's supply function is its Marginal Cost (MC) function.

MC = dC(q)/dq = 0.08q

Therefore, firm's supply function: P = 0.08q

Since there are 6 firms, market supply (Qs) = 6q

q = Qs/6

P = 0.08 x (Qs/6)

Qs = 6P/0.08

Qs = 75P

(Q2) (e)

In equilibrium, Qd = Qs.

1600 - 325P = 75P

1600 = 400P

P = $4